OpenStudy (anonymous):

y = (-3)/(x+4) find the inverse of the function

OpenStudy (anonymous):

@Nnesha

OpenStudy (xapproachesinfinity):

do you have any ideas how to start?

OpenStudy (xapproachesinfinity):

what did you do in class?

OpenStudy (anonymous):

not really

OpenStudy (xapproachesinfinity):

nothing at all

OpenStudy (xapproachesinfinity):

don't remember what you did in class?

OpenStudy (anonymous):

No I meant I don't really know how to start off, that's what I'm trying to figure out

OpenStudy (anonymous):

is there a formula for it?

OpenStudy (xapproachesinfinity):

well im asking tell me anything you learned from class

OpenStudy (xapproachesinfinity):

about this

OpenStudy (anonymous):

Oh well I dunno much about this really I just started the section about finding inverse functions

OpenStudy (xapproachesinfinity):

well what did you start with in class there is no way your teacher/prof would assign you this if nothing is done in the class

OpenStudy (xapproachesinfinity):

in order for me to help i need to know what you know and what you don't

OpenStudy (anonymous):

ok well I just got done doing something like this f(x) = x -3 and g(x) = x-5) find f(g(5))

OpenStudy (xapproachesinfinity):

Okay i will assume you don't know anything yet about inverses!

OpenStudy (anonymous):

That is what I said above

OpenStudy (anonymous):

@Nnesha

OpenStudy (xapproachesinfinity):

so what is an inverse function basically and inverse it the operation we apply to undo some operation say we have 5+(3)=8 to undo addition we take 8-3=5 i started with 5 added 3 to get outcome 8 the inverse of that is start with 8 subtract 3 to get back to 5 more formally if have y=x-3 the inverse is x=y+3 then i just do a trick and switch the variable x and y so y=x+3

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so mine would be like x = (4y - 3)/(y)?

OpenStudy (xapproachesinfinity):

no just follow...

OpenStudy (xapproachesinfinity):

now i have y=4/(x-1) so according to what i just did 1/4 y=1/(x-1)

OpenStudy (xapproachesinfinity):

then i flip the fractions i get 4/y =x-1 yes?

OpenStudy (xapproachesinfinity):

i did error in the previous one so pay attention to this one carefully

OpenStudy (xapproachesinfinity):

once we solve this one you should do the same to solve yours

OpenStudy (anonymous):

ok...

OpenStudy (xapproachesinfinity):

\(\huge y=\frac{4}{x-1}\) this is our function okay

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

first: multiply by 1/4 both sides \(\huge \frac{1}{4}y=\frac{1}{4}\frac{4}{x-1}\)

OpenStudy (xapproachesinfinity):

following...

OpenStudy (xapproachesinfinity):

so far okay?

OpenStudy (anonymous):

I think so

OpenStudy (xapproachesinfinity):

so i get \(\huge \frac{1}{4}y=\frac{1}{x-1}\) 4 canceled out (4/4=1)

OpenStudy (xapproachesinfinity):

now flip the fraction \(\huge \frac{4}{y}=\frac{x-1}{1}\)

OpenStudy (xapproachesinfinity):

good!

OpenStudy (xapproachesinfinity):

??

OpenStudy (xapproachesinfinity):

answer say no, yes say at least something

OpenStudy (anonymous):

lol I'm thinking

OpenStudy (xapproachesinfinity):

okay

OpenStudy (anonymous):

it looks okay I think...

OpenStudy (xapproachesinfinity):

make sure not just okay hehe

OpenStudy (xapproachesinfinity):

make sure you got it lol

OpenStudy (anonymous):

I think I got it

OpenStudy (xapproachesinfinity):

okay

OpenStudy (anonymous):

so for mine I have to multiply both sides by 4/3

OpenStudy (xapproachesinfinity):

x-1/1 is just x-1 dividing by 1 does not change anything so i get \(\huge \frac{4}{y}=x-1\)

OpenStudy (xapproachesinfinity):

no by -1/3 to cancel -3 on top of the left part fraction

OpenStudy (xapproachesinfinity):

why 4/3?

OpenStudy (xapproachesinfinity):

this needs some good algebra you know

OpenStudy (anonymous):

cause I thought you were flipping the denominator with the numerator and then multiplying both sides by it

OpenStudy (xapproachesinfinity):

i first multiplied to cancel then i flipped the both sides fractions so the equality stays correct

OpenStudy (xapproachesinfinity):

if i had 1/y=x/2 flipping would give me y=2/x yes

OpenStudy (xapproachesinfinity):

that's what i did to that fraction

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

okay so multiply both sides by -1/3 and then I'll get

OpenStudy (xapproachesinfinity):

back to our problem we are at \(\huge \frac{4}{y}=x-1\) now i just need to add 1 to both sides to get \(\huge \frac{4}{y}+1=x\) then i switch x and y to get \(\huge \frac{4}{x}+1=y\)

OpenStudy (anonymous):

oh ok.

OpenStudy (xapproachesinfinity):

once you multiply you cancel -3 on top you get -1/3 y=1/(x+4)

OpenStudy (xapproachesinfinity):

the one i just solved got me to \(\huge y=\frac{4}{x}+1\) this the inverse i was looking for for my case

OpenStudy (xapproachesinfinity):

you follow the same steps to get there

OpenStudy (anonymous):

*bangs head against wall* I might have missed something from the last chapter this seems so convoluted

OpenStudy (xapproachesinfinity):

well one thing you can't go straight solving problems without looking at you notes first

OpenStudy (xapproachesinfinity):

it does not work that way

OpenStudy (xapproachesinfinity):

yo need to study you notes and texbook before tackling any problem

OpenStudy (anonymous):

yea, I think I'm going back to my notes and try re-teaching myself some of this stuff and then I'll just use the stuff you've written as a reference for things

OpenStudy (anonymous):

anyrate I thank you for the work you've done. not many people are patient enough to spend that much time trying to help someone so thank you

OpenStudy (xapproachesinfinity):

yes! need me just tag @x...

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

no problem! i could have just given you the answer but that won't help at all :)

OpenStudy (xapproachesinfinity):

good luck :)